OpenStudy (anonymous):

really easy question i just cant seem to grasp :) simplify - 3^4x9^2x27^3

OpenStudy (anonymous):

-3^4 = 81 9^2 = 81 27^3 = 19683 81 x 81 x 19683 = 129,140,163

OpenStudy (anonymous):

are those x's variables or did you intend those to be multiplication?